# Early-Exit Graph Neural Networks

### v1

@@ F.3 Analysis on the use of ReLU+TanH as activation
Here, we want to understand how our instantiation of the non-linearities $\sigma_{1}(x)=\text{ReLU}(\text{TanH}(x))$ and $\sigma_{2}=\text{ReLU}(x)$ results in effective w.r.t. more common choices in the design of neural networks, such as ReLU or TanH. In order to assess these aspects, we compare SAS-GNN against A-DGN and GRAFF, using all the possible combinations of ReLU, TanH, and ReLU+TanH. Our objective is to understand whether we have any advantage in terms of performance, rather than a result that is limited to the theory. We present such a comparison in Table 10. From the Table, we see that our principled approach ranks first in 2 out of 5 datasets, namely Tolokers and Questions. While A-DGN performs significantly better than GRAFF and SAS-GNN in Minesweeper and Roman Empire. In the other datasets, the performance remains comparable for all the activation types. Surprisingly, GRAFF tends to perform worse than the other models, even though it was designed specifically to handle heterophilic graphs. However, when it was proposed by Di Giovanni et al. [2023], this specific benchmark was not introduced yet, thus, we attribute this gap to this new collection. We present additional evidence on the impact of ReLU+TanH, when used within other GNNs, such as GCN, GraphSAGE, GAT, and GIN, in Section F.6.
@@ F.4 Full Results on the benchmark for heterophily
In this section, we present the full leaderboard results from the benchmark proposed by Platonov et al. [2023]. While Table 3 already reported partial results showcasing the best-performing models, here we provide a more complete view.
@@
The benchmark from Platonov et al. [2023] focuses on evaluating GNNs on a suite of diverse, medium-to-large graphs with varying degrees of heterophily, domain, density, and size. In their study, it was demonstrated that many GNN models originally designed to handle heterophily underperform compared to classical GNNs, such as GCN, GraphSAGE, and GAT, or even GTs, when evaluated under this more realistic setting.
Given this context, our proposed SAS-GNN and EEGNN models exhibit superior performance relative to heterophily-specialized methods, and are comparable to, or even outperform, the classic GNNs in several cases. Additionally, we include results for the same classical models after undergoing a more extensive hyperparameter tuning, as reported by Luo et al. [2024]. These tuned versions incorporate techniques such as dropout, batch normalization, and residual connections, which lead to significant performance improvements for both message-passing architectures (e.g., GCN, GraphSAGE, GAT) and Graph Transformers (e.g., Polynormer).
Despite not employing normalization or dropout in our models, SAS-GNN and EEGNN remain competitive on Tolokers and Questions, where we still match or surpass several strong baselines. However, on other datasets, we acknowledge a noticeable gap when compared to extensively tuned models. This observation further highlights the importance of architectural principles that incorporate mechanisms like weight sharing and depth-resilient message-passing. This enables our models to maintain competitive performance in diverse settings without relying on auxiliary components.
@@
Table 11: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. The models marked with the asterisks are those presented in Table 3, reported by Luo et al. [2024].
@@
Classic GNNs Platonov et al. [2023]
@@
Heterophily-Specialized Models Platonov et al. [2023]
@@
Reassessment + Current SOTA Luo et al. [2024]
@@
CO-GNN 54.17 ± 0.37 97.31 ± 0.41 91.37 ± 0.35 84.45 ± 1.17 76.54 ± 0.95
Ours Luo et al. [2024]
SAS-GNNnoedge 51.47 ± 0.68 93.29 ± 0.61 83.46 ± 0.61 85.80 ± 0.79 79.60 ± 1.15
EEGNN 51.47 ± 0.51 93.18 ± 1.37 80.36 ± 0.43 85.26 ± 0.65 78.90 ± 1.15
F.5 Full Results on Long Range Graph Benchmark
In the main paper, we discussed the performance of SAS-GNN and EEGNN against MPNNs and GTs, as well as asynchronous message-passing methods such as Co-GNN and AMP. We limited our analysis, since we consider them as our main competitor on the LRGB task. However, other approaches exist that attempt to improve the GNNs’ capabilities to overcome their long-range information propagation capabilities. Among these, we have Rewiring Methods that also contribute to high scoring in long-range tasks. Here we include DIGL [Gasteiger et al., 2022], MixHop [Abu-El-Haija et al., 2019], and DRew [Gutteridge et al., 2023]. The results are in Table 12. We also distinguish here between SAS-GNNnoedge ($f_{e}(\mathbf{E})=0$), SAS-GNNedge ($f_{e}(\mathbf{E})=\mathbf{BEW}_{e}$), and SAS-GNN/EEGNNours ($f_{e}(\mathbf{E})=-\text{ReLU}(\mathbf{BEW}_{e})$). While we do not surpass all the models, we outperform classic MPNNs in Peptides-func and Peptides-struct with SAS-GNNnoedge. Notably, edge features do not enhance performance in peptide datasets, though our edge-based method slightly improves over SAS-GNNedge. However, in Pascal-VOC, classic MPNNs achieve better results, yet our edge feature strategy improves F1, highlighting the informativeness of edge features in this domain. Against rewiring methods, we remain competitive. In Peptides-func, we surpass most models except DRew-GCN and its LapPE variant. In Peptides-struct, SAS-GNNedge underperforms, but all other versions achieve superior results. While DRew-GCN excels in peptide tasks, we outperform it in Pascal-VOC. Compared to DIGL, we perform better in peptide datasets but fall behind in Pascal-VOC. As concerns GTs, we outperform them in peptide settings except in inductive node classification. Against asynchronous MPNNs, we are slightly worse than Co-GNN in Peptides-func and comparable to AMP in Peptides-struct, despite these models being more expressive than the 1-WL test. No comparisons are available for Pascal-VOC in this category.
@@ F.5 Full Results on Long Range Graph Benchmark
Table 12: Performance comparison of models across LRGB datasets. The scores are marked in red for the first, blue for the second, and green for the third.
@@
F.6 Results using Straight Trough Gumbel Softmax on Classic GNNs
@@
In this section, we demonstrate that our early-exit module, based on the Straight-Through Gumbel-Softmax Estimator, is modular and can be integrated with various GNN architectures, including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], GAT [Veličković et al., 2018], and GIN [Xu et al., 2019].
[…]
@@
All models use the same hyperparameters as the best-performing EEGNN configuration in Table 3; these configurations can be found in the code repository. It is important to note that these implementations differ from those in prior benchmarks such as Luo et al. [2024], as we apply weight sharing and operate within our specific experimental framework.
Our goal is to analyze how integrating the early-exit (EE) module affects each GNN’s performance. Table 13 summarizes this comparison. We highlight in bold all instances where using the EE module leads to performance improvements. These gains can be found more with models trained with ReLU, and are particularly pronounced when the base model struggles (e.g., GAT on Minesweeper or GIN on Questions). We hypothesize that these improvements stem from avoiding excessive depth, which otherwise leads to performance degradation due to over-processing. In such cases, the EE module enables early termination and stabilizes performance.
Notably, SAS-GNN does not suffer from such degradation, as shown earlier. Its design inherently mitigates over-smoothing and over-squashing, making it robust to deeper configurations. Conversely, the EE module can negatively impact performance in some setups—for instance, GIN on Minesweeper and GraphSAGE on Tolokers. This degradation likely arises from applying weight sharing to models not originally designed for it, potentially reducing their expressivity due to fewer parameters. Still, exceptions exist—e.g., GraphSAGE achieves unexpectedly high AUROC on Minesweeper, comparable to the values in Table 3.
Since SAS-GNN and EEGNN benefit from using a combination of ReLU and Tanh nonlinearities, as motivated in Corollary 3.3, we extended this configuration to the other GNNs. In Table 13, we mark with an asterisk all cases where ReLU+Tanh improves performance over ReLU alone. We find that this hybrid activation consistently enhances results across most settings.
To summarize, the early-exit module can significantly improve performance, particularly when the base model struggles with deeper architectures. In such cases, EE enables earlier termination and more stable outcomes. In other cases, both the base and EE-enhanced models may underperform, likely due to the limitations of weight sharing in architectures not originally designed for it. Nevertheless, we find evidence—such as GraphSAGE on Minesweeper—that weight sharing can still be effective on certain datasets. Additionally, using ReLU+Tanh, as proposed in SAS-GNN, proves beneficial even for classical GNNs, improving performance and narrowing the gap between base and EE-enhanced models.
@@
Table 13: Performance of classic GNNs using Gumbel Softmax as early-exit. Models are also tested by changing the activation function type.
@@
SAS-GNN - $93.29\pm 0.61$ $79.60\pm 1.15$ $85.80\pm 0.79$
EEGNN - $93.18\pm 1.37$ $78.90\pm 1.15$ $85.26\pm 0.65$
F.7 Results on Homophilic Node Classification

### v2

@@ F.3 Analysis on the use of ReLU+TanH as activation
Here, we want to understand how our instantiation of the non-linearities $\sigma_{1}(x)=\text{ReLU}(\text{TanH}(x))$ and $\sigma_{2}=\text{ReLU}(x)$ results in effective w.r.t. more common choices in the design of neural networks, such as ReLU or TanH. In order to assess these aspects, we compare SAS-GNN against A-DGN and GRAFF, using all the possible combinations of ReLU, TanH, and ReLU+TanH. Our objective is to understand whether we have any advantage in terms of performance, rather than a result that is limited to the theory. We present such a comparison in Table 16. From the Table, we see that our principled approach ranks first in 2 out of 5 datasets, namely Tolokers and Questions. While A-DGN performs significantly better than GRAFF and SAS-GNN in Minesweeper and Roman Empire. In the other datasets, the performance remains comparable for all the activation types. Surprisingly, GRAFF tends to perform worse than the other models, even though it was designed specifically to handle heterophilic graphs. However, when it was proposed by [Di Giovanni et al., 2023], this specific benchmark was not introduced yet, thus, we attribute this gap to this new collection. We present additional evidence on the impact of ReLU+TanH, when used within other GNNs, such as GCN, GraphSAGE, GAT, and GIN, in Section F.6.
@@ F.4 Full Results on the Benchmark for Heterophily
In this section, we present the full leaderboard results from the heterophily benchmark introduced by Platonov et al. [2023]. While Table 3 already provided a fair comparison with our approach, here we aim to give a more complete picture.
@@
The benchmark of Platonov et al. [2023] evaluates GNNs on diverse medium-to-large graphs spanning different domains and exhibiting varying degrees of heterophily, density, and size. Their study revealed that many models explicitly designed for heterophily often underperform compared to classical GNNs such as GCN, GraphSAGE, GAT, or even GTs when tested under this more realistic setting. More recently, Luo et al. [2024] showed that the performance of these classic MPNNs can be pushed even further by incorporating simple architectural components, most notably, dropout layers interleaved by message-passing layers. Table 17 reports the updated benchmark. Results are taken by Platonov et al. [2023], Finkelshtein et al. [2024], Luo et al. [2024].
From this broader context, two main observations emerge with respect to our approach. First, our models consistently outperform heterophily-specialized architectures, even without extensive tuning, demonstrating that our design generalizes well to heterophilic settings. Second, performance gains in these benchmarks are strongly correlated with components that are not currently supported by our theoretical framework, such as dropout or normalization layers. Indeed, ablating dropout leads to significant performance drops across baselines. In some cases (e.g., Minesweeper and Roman Empire) models without dropout still surpass ours, while in others (e.g., Questions and Tolokers) dropout is less critical and our models perform comparably or even better. These cases highlight situations where our principled design is particularly effective.
At the same time, these results underscore an important limitation of principled architectures: in practical scenarios, accuracy improvements often require the inclusion of black-box components that fall outside current theoretical guarantees. This gap between theoretical soundness and empirical performance remains a key challenge for future work. We included in the table also some non-principled version of SAS-GNN, with no hyperparameter tuning, where we remove weight-sharing, or include dropout, or include normalization, and all the possible combination of these components. Here we see, that with no hyperparameter tuning the test performance already improves significantly. An example of this is Roman Empire ($83.46\rightarrow 87.32$). We believe that our model could benefit from an extensive hyperparameter tuning, however this questions its capabilities for safe early exits.
Neural Adaptive Step. We further analyze the role of our Neural Adaptive Step mechanism in node classification, an aspect not covered in the main paper. Specifically, we compare the default formulation, where each node is assigned its own adaptive step vector, against a simplified variant where the step is reduced to a scalar shared across all nodes. To ensure differentiability, the scalar is defined as the mean of the vector entries. Concretely, in Algorithm 1, the update rule becomes:
$\mathbf{H}^{l+1}\leftarrow\mathbf{H}^{l}+\mathrm{mean}(\bm{\tau}^{l})\,\Delta\mathbf{H}^{l},$
so that every node is updated with the same constant factor, namely the average prediction of the confidence network.
This experiment highlights the impact of personalized integration constants. While the scalar variant enforces a uniform update across all nodes, the full neural adaptive step allows each node to adapt its integration constant based on its own confidence estimate. As shown in Table 17, moving from the vector to the scalar consistently degrades performance across datasets, with particularly pronounced drops on Questions. We attribute this to the fact that averaging removes much of the information conveyed by the confidence network, weakening its guidance to the message-passing process.
We do not extend this analysis to graph-level tasks, as in those cases, we already use a scalar integration constant by design, since there the confidence prediction already operates at the graph level. Nevertheless, these results provide strong evidence that node-wise adaptive steps are essential for capturing heterogeneity in node-level tasks, and that personalized integration constants substantially contribute to the effectiveness of our approach.
@@
Table 17: Node classification under heterophily. The scores are marked in red for the first, blue for the second, and green for the third. The models marked with the asterisks are those presented in Table 3, reported by Luo et al. [2024]. Results are taken by Platonov et al. [2023], Deng et al. [2024], Finkelshtein et al. [2024], Luo et al. [2024].
@@
Classic GNNs [Platonov et al., 2023]
@@
Heterophily-Specialized Models [Platonov et al., 2023]
@@
Reassessment + Current SOTA [Luo et al., 2024]
@@
CO-GNN [Finkelshtein et al., 2024] 54.17 ± 0.37 97.31 ± 0.41 91.37 ± 0.35 84.45 ± 1.17 76.54 ± 0.95
Ours
SAS-GNN $51.47\pm 0.68$ $93.29\pm 0.61$ $83.46\pm 0.61$ $85.80\pm 0.79$ $79.60\pm 1.15$
SAS-GNN-ws $51.51\pm 0.41$ $92.62\pm 0.45$ $84.45\pm 0.44$ $85.46\pm 0.63$ $78.79\pm 0.97$
SAS-GNN+drop+norm $51.19\pm 0.53$ $93.40\pm 0.48$ $86.00\pm 0.53$ $85.98\pm 0.55$ $76.64\pm 1.40$
SAS-GNN+drop+norm-ws $52.16\pm 0.50$ $93.21\pm 0.40$ $87.32\pm 0.00$ $85.87\pm 0.70$ $77.57\pm 1.10$
EEGNN $51.47\pm 0.51$ $93.18\pm 1.37$ $80.36\pm 0.43$ $85.26\pm 0.65$ $78.90\pm 1.15$
[…]
F.5 Full Results on the Long Range Graph Benchmark
In the main paper, we compared SAS-GNN and EEGNN against classical MPNNs, Graph Transformers (GTs), asynchronous message-passing models such as Co-GNN, AMP, MTGCN [Pei et al., 2024], and graph neural ODEs, which are the class of models most closely related to ours. Among these we considered a diffusion-inspired message-passing, namely GRAND [Chamberlain et al., 2021], a physics-inspired scheme GraphCON [Rusch et al., 2022], and two examples of stable and non-dissipative methods, namely A-DGN and SWAN [Gravina et al., 2025]. SWAN is an example of global as well as local stable and non-dissipative message-passing, making it a better alternative w.r.t. A-DGN. Our models, SAS-GNN and EEGNN are limited by local non-dissipativeness.
Here, we provide a broader view of the Long Range Graph Benchmark (LRGB).
As discussed in the main text, peptide datasets are no longer considered representative of long-range dependencies [Bamberger et al., 2025], a point we also validated in Figure 4. Beyond this, we also include the revised results of Luo et al. [2025], who revisited inductive tasks and introduced GCN+, a variant that view classical GCN layers as a combination of message-passing with dropout, normalization, residual connections, and feedforward layers, inspired by Transformer architectures. In practice, dropout and normalization were the most influential components. We further include rewiring methods specifically designed to address topological bottlenecks [Arnaiz-Rodriguez and Errica, 2025], such as DIGL [Gasteiger et al., 2022], MixHop [Abu-El-Haija et al., 2019], and DRew [Gutteridge et al., 2023]. Results are taken from Finkelshtein et al. [2024], Gravina et al. [2025], Luo et al. [2025], Tönshoff et al. [2023], and are summarized in Table 18.
We report three SAS-GNN variants: SAS-GNNnoedge ($f_{e}(\mathbf{E})=0$), SAS-GNNedge ($f_{e}(\mathbf{E})=\mathbf{BEW}_{e}$), and SAS-GNN/EEGNNours ($f_{e}(\mathbf{E})=-\text{ReLU}(\mathbf{BEW}_{e})$). While we do not outperform all baselines, we surpass classical MPNNs on Peptides-func and Peptides-struct with SAS-GNNnoedge. Interestingly, edge features offer little benefit on peptides, though our edge-based formulation still slightly improves over SAS-GNNedge. On Pascal-VOC, however, classic MPNNs remain stronger overall, but edge features boost our F1 scores, showing their usefulness in this domain.
Against rewiring methods, our models remain competitive: in Peptides-func, we outperform most approaches except DRew-GCN and its LapPE variant; in Peptides-struct, SAS-GNNedge underperforms, but the other variants remain strong. While DRew-GCN excels on peptide datasets, we surpass it on Pascal-VOC. Compared to DIGL, our methods are stronger on peptides but weaker on Pascal-VOC.
When comparing to GTs, we generally perform better in peptide tasks, except for inductive node classification. Relative to asynchronous MPNNs, we fall slightly behind Co-GNN on Peptides-func and remain comparable to AMP on Peptides-struct, despite their higher expressivity beyond 1-WL. Results for Pascal-VOC are not available in this category.
Finally, when comparing with GCN+, we observe that its performance boost in peptides primarily comes from dropout, whose effect is much smaller in our framework. Yet, as already noted, peptides are not long-range benchmarks, and no class of models, including Transformers, rewiring methods, or asynchronous approaches—shows a consistent advantage there. In this setting, black-box components appear to dominate. By contrast, Pascal-VOC constitutes a true long-range task, as evidenced by the performance gap between Transformers and MPNNs and validated in Bamberger et al. [2025]. Interestingly, even methods explicitly designed for long-range reasoning, such as A-DGN and SWAN, fall behind on this dataset, even trailing GCN+. Since ablations reveal that GCN+ benefits primarily from normalization layers, also present in Transformers, we attribute this gap not to an inability to capture long-range dependencies, but rather to the increased complexity required in message passing.
Overall, these findings suggest that while Graph Neural ODEs provide transparency and theoretical grounding, they may underperform in practical long-range scenarios where architectural heuristics (e.g., dropout, normalization) play a decisive role.
@@ F.5 Full Results on the Long Range Graph Benchmark
Table 18: Performance comparison of models across LRGB datasets. The scores are marked in red for the first, blue for the second, and green for the third. Results are taken from Finkelshtein et al. [2024], Luo et al. [2025], Gravina et al. [2025], Errica et al. [2023], Tönshoff et al. [2023].
@@
Reassessment of MPNNs [Luo et al., 2025]
GCN+ 72.61 ± 0.67 0.2421±0.0016 33.57 ± 0.87
– Dropout 67.48 ± 0.55 0.2549 ± 0.0025 30.72 ± 0.69
– Norm 71.07 ± 0.27 0.2509 ± 0.0025 18.02 ± 1.11
GatedGCN+ 70.06 ± 0.33 0.2431 ± 0.0020 42.63 ± 0.57
– Dropout 66.95 ± 1.01 0.2508 ± 0.0014 33.89 ± 0.66
– Norm 67.33 ± 0.26 0.2474±0.0015 36.28 ± 0.43
@@
MTGCN 69.36 ± 0.89 0.2461 ± 0.0019 -
Graph Neural ODEs
GRAND 57.89 ± 0.62 0.3418 ± 0.0015 19.18 ± 0.97
GraphCON 60.22 ± 0.68 0.2778 ± 0.0018 21.08 ± 0.91
A-DGN 59.75 ± 0.44 0.2874 ± 0.0021 23.49 ± 0.54
SWAN 63.13 ± 0.46 0.2571 ± 0.0018 27.96 ± 0.48
@@
F.6 Results using Straight-Through Gumbel-Softmax on Classic GNNs
@@
In this section, we demonstrate that our early-exit (EE) module, based on the Straight-Through Gumbel-Softmax estimator, is modular and can be integrated with various GNN architectures, including GCN [Kipf and Welling, 2017], GraphSAGE [Hamilton et al., 2017], GAT [Veličković et al., 2018], and GIN [Xu et al., 2019]. We first show how these models perform using the deep regime $L=20$. Then we perform a sensitivity analysis study, where we include also EEGNN, where we allow $L$ to vary across an interval, to understand how the number of parameters, exit distributions, and performance change. We already provided an example with the performance variation in the main text in Figure 3.
Deep Regime experiment. To integrate the EE module, we focus on message-passing schemes with weight sharing, which allows the shared functions $f_{c}$ and $f_{\nu}$ to operate across layers. We consider a deep regime with $L=20$, enabling each node to leverage all shared weight matrices and allowing us to study the effect of deep processing on classic MPNNs. Concretely, we replace line 1 in Algorithm 1 with the message-passing function of the desired GNN. We denote these models as EE+GNN-type, where GNN-type $\in\{\text{GCN, SAGE, GAT, GIN}\}$.
@@
All models use the same hyperparameters as the best-performing EEGNN configuration in Table 3, available in our code repository. We do not perform additional tuning, as our goal is to isolate the impact of the Neural Adaptive Step and end-to-end early-exit optimization in deep regimes. Note that these implementations differ from prior benchmarks [Luo et al., 2024], as they are evaluated within our experimental framework.
We report results on the heterophilic graph benchmarks from Platonov et al. [2023], which vary widely in density, size, and context, providing a more representative evaluation not biased by homophily assumptions. Table 19 shows results on Minesweeper, Questions, and Tolokers, while Table 20 shows results on Amazon Ratings and Roman Empire. Bold indicates EE variants outperforming their base model, and asterisks mark cases where ReLU+Tanh generally improves over ReLU.
Across the tables, we observe that EE modules generally improve performance, especially with ReLU activations. Base models often struggle in deep regimes (e.g., GAT on Minesweeper, GIN on Questions or Roman Empire), likely due to over-processing in deep layers. In these cases, EE enables early termination, stabilizing performance. SAS-GNN, by contrast, naturally avoids over-smoothing and over-squashing, making it robust to deep configurations.
Some exceptions occur where EE slightly reduces performance, such as GIN on Minesweeper or GraphSAGE on Tolokers. This may result from applying weight sharing to architectures not originally designed for it, reducing expressivity. Generally, we do not expect EE to always increase performance, as the application of Neural Adaptive Step and exiting before is require only to stabilize the performance, it does not add additional information that potential can improve performance. We also observe from these experiments that GraphSAGE generally tolerates deep processing well, showing notable improvements with EE. Additionally, ReLU+Tanh activations consistently yield stable performance across both base and EE models, suggesting this combination effectively stabilizes layer-wise feature evolution in practice. However, we still suggest SAS-GNN or the use of more principled design as a backbone to perform early-exiting, since these are supported by theoretical proofs, and remain a safer solution to perform deep processing when needed.
Instability of Early-Exit. As discussed in Section 4, the effectiveness of an early-exit head depends on the stability of node embeddings throughout their feature evolution. But what happens when deeper processing is required? In Figure 3, we illustrate this issue on the Questions dataset: early-exit performance appears sensitive to changes in the budget $L$, which is undesirable.
To further investigate, we compare EEGNN against baselines and provide additional evidence of its advantages. Figure 13 reports both (i) parameter counts (without weight sharing for classic MPNNs, and with for EEGNN) and (ii) exit distributions across different budgets $L$. On the right, we observe that weight sharing keep the number of parameters for EEGNN self-contained. On the left, EEGNN uniquely shifts its exits toward higher depths, indicating that the model attempts to fully utilize its capacity. Although we cannot directly link this behavior to problem-radius prediction, since even with knowledge of the radius, it remains unclear whether sufficient feature extraction occurs within that many layers, we can conclude that EEGNN manage also to increase its effective depth on average when adopting a training with higher budget. Consistent with this, its performance improves at $L=20$, confirming the benefit of deeper processing.
However, this trend does not always hold. For example, in the Tolokers dataset (Figure 14), we repeat the same analysis over a restricted interval of $L$ values. Here, EEGNN again shows superior stability and reduced sensitivity to budget variation. At test time, the average exit shifts toward the maximum budget, suggesting that the model tends to fully exploit the available depth. In these case, all the baselines tend to use full model capacity, but without improving the overall performance, and sometimes degrading it. We don’t want to affirm that EEGNN always improves, as $L$ increases, but rather show how it never degrades significantly. In this setting, efficiency is not the main objective; rather, the focus is on verifying the model’s ability to process up to a large $L$. Finally, thanks to weight sharing, the parameter count remains constant, enabling improved performance without increasing memory complexity.
@@
Table 19: Performance of classic GNNs using Gumbel Softmax as early-exit on Minesweeper, Questions, Tolokers. Models are also tested by changing the activation function type.
@@
SAS-GNN - $93.29\pm 0.61$ - $79.60\pm 1.15$ - $85.80\pm 0.79$
EEGNN - $93.18\pm 1.37$ - $78.90\pm 1.15$ - $85.26\pm 0.65$
Table 20: Performance of classic GNNs and SAS/EEGNN on Amazon Ratings and Roman Empire datasets. Bold indicates EE variants outperforming the base model.
Model Amazon Ratings Roman Empire
ReLU ReLU + Tanh ReLU ReLU + Tanh
GCN 49.72 $\pm$ 0.3 48.53 $\pm$ 0.4 81.16 $\pm$ 0.6 76.75 $\pm$ 0.4
EE+GCN 51.13 $\pm$ 0.7 *51.25 $\pm$ 0.4 79.89 $\pm$ 0.7 79.76 $\pm$ 0.4
GAT 48.40 $\pm$ 0.5 *49.08 $\pm$ 0.7 42.26 $\pm$ 32.4 *75.86 $\pm$ 1.1
EE+GAT 51.88 $\pm$ 0.5 51.85 $\pm$ 0.5 77.21 $\pm$ 0.9 *77.48 $\pm$ 0.8
SAGE 52.06 $\pm$ 0.4 *52.26 $\pm$ 0.6 84.28 $\pm$ 0.8 *86.55 $\pm$ 0.6
EE+SAGE 50.51 $\pm$ 0.7 *50.67 $\pm$ 0.6 78.94 $\pm$ 6.4 *81.10 $\pm$ 1.2
GIN 42.69 $\pm$ 4.3 *49.50 $\pm$ 0.5 13.96 $\pm$ 0.0 *83.27 $\pm$ 1.1
EE+GIN 51.12 $\pm$ 0.8 *51.26 $\pm$ 0.5 65.85 $\pm$ 0.8 65.75 $\pm$ 0.8
SAS-GNN - 51.47 $\pm$ 0.7 - 83.46 $\pm$ 0.6
EEGNN - 51.54 $\pm$ 0.5 - 80.36 $\pm$ 0.4
Figure 13: Exit distributions of EEGNN against the others (left) and Parameter evolution based on $L$ (right).
Figure 14: Performance variation based on budget $L$. Exit distributions of EEGNN against the others (right). Number of Parameters evolution based on $L$ (bottom).
F.7 Additional Results on Accuracy–Efficiency Trade-Off Analysis
In this section, we extend our trade-off analysis to further compare EEGNN and SAS-GNN against classical MPNNs, Graph Transformers (GTs), and asynchronous MPNNs. While our models are not always state-of-the-art in terms of raw accuracy, partly due to the stronger regularization and normalization schemes employed by competing methods, we argue that they provide clear benefits in terms of space and time efficiency.
This analysis differs from the runtime comparison in Table 2, as here we evaluate models under their best-performing hyperparameter configurations, namely the hidden dimension $m^{\prime}$ and number of layers $L$ (summarized in Table 21). In the main paper, we showed that SAS-GNN and EEGNN achieve competitive accuracy on several benchmarks. Here, we complement those results with experiments on Amazon Ratings and Roman Empire, two datasets where our models underperform relative to GTs or heavily regularized MPNNs.
The results, illustrated in Figures 1516, highlight the main advantage of our approach. Even when not achieving the highest accuracy, SAS-GNN and EEGNN consistently deliver superior space efficiency, requiring significantly fewer parameters. In terms of inference time, our models are also among the fastest across both datasets, with the sole exception of GCN on Roman Empire. These findings reinforce our claim that EEGNN and SAS-GNN strike a favorable balance between accuracy, scalability, and efficiency.
(a) Accuracy vs Parameters
(b) Accuracy vs Inference Time
Figure 15: Trade-off analysis on Amazon Ratings.
(a) Accuracy vs Parameters
(b) Accuracy vs Inference Time
Figure 16: Trade-off analysis on Roman Empire.
Table 21: Model configurations across Questions, Roman Empire, and Amazon Ratings during trade-off analysis. We report hidden dimension $m^{\prime}$ and number of layers $L$.
Model Questions $(m^{\prime},L)$ Roman Empire $(m^{\prime},L)$ Amazon Ratings $(m^{\prime},L)$
SAS-GNN (64, 9) (128, 12) (256, 9)
EEGNN (64, 20) (128, 20) (256, 20)
Co-GNN (64, 9) (512, 12) (256, 9)
PolyNormer (64, 10) (64, 20) (256, 20)
SGFormer (64, 10) (256, 10) (64, 10)
NodeFormer (64, 10) (256, 10) (64, 10)
GCN (512, 10) (256, 9) (512, 4)
F.8 Additional Results on APGCN vs. EEGNN
As we have analysed and commented in the main text, the advantage of our EEGNN is that it correctly update the exit module based on the task, and this gives a potential to identify the correct exit point, as well as preserving feature informativeness when it exits. We now analyse such a behavior on the Eccentricity dataset, where a similar trend is observed. As illustrated in Figure 17, APGCN fails to propagate information sufficiently, exiting early with an average depth of only $7.10\pm 0.74$, resulting in a high Test MAE of $11.08$. In contrast, EEGNN correctly identifies the complexity of the task, adapting to a significantly deeper regime with an average of $16.71\pm 6.76$ steps, an average value even higher than that observed for SSSP. While the performance is not absolute state-of-the-art, the error is reduced by over 50% compared to APGCN (MAE $5.05$). This confirms that when deep processing is required, but we don’t want to manually tune the depth it is crucial to learn how to select the exit point, and having stable intermediate representations. EEGNN leverages these representations to sustain deeper propagation without succumbing to the premature exiting behavior typical of budget-constrained mechanisms.
Figure 17: Distribution of exit steps on the Eccentricity dataset over 4 seeds. Left: APGCN exits prematurely (Mean $\approx 7$) leading to high error. Right: EEGNN adapts to the task’s long-range requirements by sustaining deeper propagation (Mean $\approx 17$), significantly reducing the Test MAE.
F.9 Results on Homophilic Node Classification